OpenStudy (anonymous):

Solve the inequality (x/2) > (-8)/(x + 4) + 3 for domain

OpenStudy (anonymous):

Multiply through with 2(x+4)

OpenStudy (anonymous):

Soooo x > ( -16/2x +8) + 6?

OpenStudy (anonymous):

You are way off. The first term should start out like this: \[(x(2)(x+2) \div 2)\] If you notice the 2s cancel.

OpenStudy (anonymous):

Im sorry im really confused as to whats going on...

OpenStudy (anonymous):

Im sorry im really confused as to whats going on...

OpenStudy (anonymous):

The terms are in the form of fractions. In order to make it simpler you can get rid of the fractions. In order to get rid of the fractions, you multiply each term by 2(x+4)

OpenStudy (anonymous):

Just to make sure were looking at the same thing, the original problem looks like this \[x \div2 > \left( \left( -8 \right)\div \left( x+4 \right) \right) + 3\]

OpenStudy (anonymous):

Yes pay attention to the terms in the bottom (denominator): 2 and (x + 4). You have to get rid of the fractions to make it easier to handle.

OpenStudy (anonymous):

sooo its \[x \times 2(x + 4) > -8 \times 2(x + 4) + 3\times 2(x + 4)\]

OpenStudy (anonymous):

You are seeing OK but you are not performing the operations. May be write it out on a piece of paper. You can not just make the bottom (denominator) disappear by magic. You have to cancel it out with something. You cross out the 2 at the bottom AND cross out the 2 on top. On the other side cross out the (x+4) on the bottom AND (x+4) on top. Your answer would be slightly different. But you are beginning to see it.